# Actions Speak Louder than Words: Trillion-Parameter Sequential Transducers for Generative Recommendations

### v1

@@ 4.1.1 Traditional Sequential Settings
We first evaluate the performance of HSTU on two popular recommender datasets, MovieLens and Amazon Reviews. We follow sequential recommendation settings in literature, including full shuffle and multi-epoch training. For baseline, we use SASRec, a state-of-the-art Transformer implementation (Kang & McAuley, 2018). We report Hit Rate@K and NDCG@K over the entire corpus, consistent with recent work (Dallmann et al., 2021; Zhai et al., 2023a).
@@ 4.1.2 Industrial-scale Streaming Settings
We show results in Table 4. First, HSTU significantly outperforms Transformers, especially in ranking, likely due to pointwise attention and improved relative attention biases. Second, the gaps between the ablated HSTUs and HSTU confirm the effectiveness of our designs. Optimal learning rates are about 10x lower for Softmax-based HSTU and Transformer vs the rest due to training stability. Even with lower learning rates and pre-norm residual connections (Xiong et al., 2020), we encountered frequent loss explosions with standard Transformers in ranking. Finally, HSTU outperforms a popular Transformer variant used in LLMs, Transformer++ (Touvron et al., 2023a), which uses RoPE (Su et al., 2023), SwiGLU, etc. Overall, in this small-scale setting, HSTU shows better quality at 1.5x-2x faster wall-clock time and uses 50% less HBM usage.
@@ 4.2 Encoder Efficiency
Stochastic Length. Figure 4 and Figure 5 (a) show the impact of stochastic length (SL) on model metrics. At $\alpha=1.6$, a sequence of length $4096$ is turned into a sequence of length $776$ the majority of the time, or removing more than 80% of the tokens. Even after sparsity ratio increases to 64%-84%, the NEs we obtained for main tasks did not degrade by more than 0.002 (0.2%). This evidence supports that SL, for suitable $\alpha$s, does not negatively impact model quality and allows for high sparsity to reduce training cost. We further verify in Section D.3 that SL significantly outperforms existing length extrapolation techniques.
Figure 4: Impact of Stochastic Length (SL) on metrics. Left: $n=4096$. Right: $n=8192$. Full results can be found in Appendix D.
@@ 4.3 Generative Recommenders vs DLRMs in Industrial-scale Streaming Settings
Lastly, we compare the end-to-end performance of GRs against state-of-the-art DLRM baselines in industrial-scale streaming settings. Our GR implementation reflects a typical configuration used in production, whereas the DLRM configurations reflect continued iterations of hundreds of people over multiple years. Given multiple generators are generally used in the retrieval stage of a recommendation system, we report both the online result for adding GR (“add source”) and replacing existing main DLRM source (“replace source”).  Table 5 and Table 6 show that GR not only significantly outperforms DLRMs offline, but also brings 12.4% wins in online A/B tests.
@@
Figure 6: Comparison of inference throughput, in the most challenging ranking setup. Full results can be found in Appendix E.
We finally compare the efficiency of GRs with our production DLRMs in Figure 6. Despite the GR model being 285x more complex in terms of FLOPs, we achieved 1.50x/2.48x higher throughput when scoring 1024/16384 candidates, respectively, due to the efficient HSTU encoder design and the M-FALCON serving algorithm presented in Section 3.4.
@@ 4.3.1 Scaling Law for Recommendation Systems
Results are shown in Figure 7. In the low compute regime, DLRMs might outperform GRs due to handcrafted features, corroborating the importance of feature engineering in traditional DLRMs. However, GRs demonstrate substantially better scalability with respect to FLOPs when continuing to increase the model complexity, whereas DLRM performance plateaus, consistent with findings in prior work. We also observe better scalability w.r.t. both embedding parameters and non-embedding parameters, with GRs leading to 1.5 trillion parameter models, whereas DLRMs performance saturate at about 200 billion parameters.
Finally, all of our main metrics, including Hit Rate@100 and Hit Rate@500 for retrieval, and NE for ranking, empirically scale as a power law of compute used given appropriate hyperparameters. We observe this phenomenon across three orders of magnitude, up till the largest models we were able to test (8,192 sequence length, 1,024 embedding dimension, 24 layers of HSTU), at which point the total amount of compute we used (normalized over 365 days as we use a standard streaming training setting) is close to the total training compute used by GPT-3 (Brown et al., 2020) and LLaMa-2 (Touvron et al., 2023b), as shown in Figure 1. Within a reasonable range, the exact model hyperparameters play less important roles compared to the total amount of training compute applied. In contrast to language modeling (Kaplan et al., 2020), sequence length play a significantly more important role in GRs, and it’s important to scale up sequence length and other parameters (e.g., embedding dimension) in tandem. This is perhaps the most important advantage of our proposed method, as we’ve shown for the first time that scaling law from LLMs (Kaplan et al., 2020) may also apply to large-scale recommendation systems.
@@ 5 Related Work
Prior work on sequential recommenders assumes that user interactions can be reduced to a single homogeneous sequence over items (Kang & McAuley, 2018; Sun et al., 2019). Industrial-scale applications of sequential approaches are primarily pairwise attention (Zhou et al., 2018) or sequential encoders as part of DLRMs (Chen et al., 2019; Xia et al., 2023). Multi-stage attention has been explored in lieu of self-attention to improve efficiency (Chang et al., 2023). Generative approaches that represent ids as a token series have been explored in retrieval (Zhuo et al., 2020).
@@ Acknowledgements
This work represents the joint efforts of hundreds of people, and would not be possible without work from the following contributors (alphabetical order): Adnan Akhundov, Bugra Akyildiz, Shabab Ayub, Alex Bao, Renqin Cai, Jennifer Cao, Guoqiang Jerry Chen, Lei Chen, Sean Chen, Huihui Cheng, Weiwei Chu, Ted Cui, Shiyan Deng, Nimit Desai, Fei Ding, Francois Fagan, Lu Fang, Liang Guo, Liz Guo, Jeevan Gyawali, Yuchen Hao, Daisy Shi He, Jie Hua, Yanzun Huang, Hongyi Jia, Rui Jian, Jian Jin, Rahul Kindi, Changkyu Kim, Fu Li, Hong Li, Shen Li, Wei Li, Zhijing Li, Xueting Liao, Emma Lin, Hao Lin, Jingzhou Liu, Xingyu Liu, Kai Londenberg, Liang Luo, Linjian Ma, Matt Ma, Yun Mao, Bert Maher, Matthew Murphy, Satish Nadathur, Min Ni, Jongsoo Park, Jing Qian, Lijing Qin, Alex Singh, Timothy Shi, Dennis van der Staay, Xiao Sun, Colin Taylor, Shin-Yeh Tsai, Rohan Varma, Omkar Vichare, Alyssa Wang, Pengchao Wang, Shengzhi Wang, Wenting Wang, Xiaolong Wang, Zhiyong Wang, Wei Wei, Bin Wen, Eric Xu, Bi Xue, Zheng Yan, Chao Yang, Junjie Yang, Zimeng Yang, Chunxing Yin, Daniel Yin, Yiling You, Keke Zhai, Yanli Zhao, Zhuoran Zhao, Hui Zhang, Jingjing Zhang, Lu Zhang, Lujia Zhang, Na Zhang, Rui Zhang, Xiong Zhang, Ying Zhang, Zhiyun Zhang, Charles Zheng, Erheng Zhong, Xin Zhuang. We would like to thank Shikha Kapoor, Rex Cheung, Lana Dam, Ram Ramanathan, Nipun Mathur, Bo Feng, Yanhong Wu, Zhaohui Guo, Hongjie Bai, Wen-Yun Yang, Zellux Wang, Arun Singh, Bruce Deng, Yisong Song, Haotian Wu, Meihong Wang for product support, and Joseph Laria, Akshay Hegde, Abha Jain, Raj Ganapathy for assistance with program management. Finally, we would like to thank Ajit Mathews, Shilin Ding, Hong Yan, Lars Backstrom for their leadership support, and insightful discussions with Andrew Tulloch, Liang Xiong, Kaushik Veeraraghavan, and Gaofeng Zhao.
@@ Appendix A Notations
We summarize key notations used in this paper in Table 7.
@@
$\Phi_{i}(t_{j})$ The $i$-th training example emitted at time $j$.
$n$ ($n_{i}$) Number of tokens in user history (of user/sample $i$).
$s_{u}(n_{i})$, $\hat{s}_{u}(n_{i})$ Sampling rate for user $i$.
$N$ $\max_{i}n_{i}$
$x_{0},\ldots,x_{n-1}$ List of input tokens.
$y_{0},\ldots,y_{n-1}$ List of output tokens.
@@
$a_{0},\ldots,a_{n-1}$ List of user actions (e.g., like, image view).
$m_{0},\ldots,m_{n-1}$ List of binary masks that we apply to $y_{0},\ldots,y_{n-1}$.
$\mathbb{X}$, $\mathbb{X}_{c}$ Vocabulary of all input/output tokens and its content subset.
@@
$d_{qk}$ Attention dimension size in HSTU and Transformers.
$d_{ff}$ Hidden dimension size in feedforward layers of Transformers.
@@
$\alpha$ Hyperparameter controlling sparsity in the Stochastic Length algorithm used in HSTU.
$\Gamma(\cdot,L)$ A function that selects a subsequence of length $L$ in Stochastic Length (SL).
$b_{m}$ Microbatch size, in the M-FALCON algorithm.
$R$ Register size.
$E,F,G,H$ Categorical features in Figure 2 (see (Mudigere et al., 2022)).
Table 7: Table of Notations
Appendix B Synthetic Data

### v2

@@ 4.1.1 Traditional Sequential Settings
We first evaluate the performance of HSTU on two popular recommender datasets, MovieLens and Amazon Reviews. We follow sequential recommendation settings in literature, including full shuffle and multi-epoch training. For baseline, we use SASRec, a state-of-the-art Transformer implementation (Kang & McAuley, 2018)  [^1] Results for other baselines are reported in Appendix D.. We report Hit Rate@K and NDCG@K over the entire corpus, consistent with recent work (Dallmann et al., 2021; Zhai et al., 2023a).
@@ 4.1.2 Industrial-scale Streaming Settings
We show results in Table 5. First, HSTU significantly outperforms Transformers, especially in ranking, likely due to pointwise attention and improved relative attention biases. Second, the gaps between the ablated HSTUs and HSTU confirm the effectiveness of our designs. Optimal learning rates are about 10x lower for Softmax-based HSTU and Transformer vs the rest due to training stability. Even with lower learning rates and pre-norm residual connections (Xiong et al., 2020), we encountered frequent loss explosions with standard Transformers in ranking. Finally, HSTU outperforms a popular Transformer variant used in LLMs, Transformer++ (Touvron et al., 2023a), which uses RoPE (Su et al., 2023), SwiGLU, etc. Overall, in this small-scale setting, HSTU shows better quality at 1.5x-2x faster wall-clock time and 50% less HBM usage.
@@ 4.2 Encoder Efficiency
Stochastic Length. Figure 4 and Figure 5 (a) show the impact of stochastic length (SL) on model metrics. At $\alpha=1.6$, a sequence of length $4096$ is turned into a sequence of length $776$ the majority of the time, or removing more than 80% of the tokens. Even after sparsity ratio increases to 64%-84%, the NEs we obtained for main tasks did not degrade by more than 0.002 (0.2%). This evidence supports that SL, for suitable $\alpha$s, does not negatively impact model quality and allows for high sparsity to reduce training cost. We further verify in Section F.3 that SL significantly outperforms existing length extrapolation techniques.
Refer to captionRefer to caption
Figure 4: Impact of Stochastic Length (SL) on metrics. Left: $n=4096$. Right: $n=8192$. Full results can be found in Appendix F.
@@ 4.3 Generative Recommenders vs DLRMs in Industrial-scale Streaming Settings
Lastly, we compare the end-to-end performance of GRs against state-of-the-art DLRM baselines in industrial-scale streaming settings. Our GR implementation reflects a typical configuration used in production, whereas the DLRM settings reflect iterations of hundreds of people over multiple years. Given multiple generators are used in the retrieval stage of a recommendation system, we report both the online result for adding GR (“add source”) and replacing existing main DLRM source (“replace source”).  Table 6 and Table 7 show that GR not only significantly outperforms DLRMs offline, but also brings 12.4% wins in A/B tests.
@@
Figure 6: Comparison of inference throughput, in the most challenging ranking setup. Full results can be found in Section H.1.
We finally compare the efficiency of GRs with our production DLRMs in Figure 6. Despite the GR model being 285x more computationally complex, we achieved 1.50x/2.99x higher QPS when scoring 1024/16384 candidates, due to HSTU and the novel M-FALCON algorithm in Section 3.4.
@@ 4.3.1 Scaling Law for Recommendation Systems
Results are shown in Figure 7. In the low compute regime, DLRMs might outperform GRs due to handcrafted features, corroborating the importance of feature engineering in traditional DLRMs. However, GRs demonstrate substantially better scalability with respect to FLOPs, whereas DLRM performance plateaus, consistent with findings in prior work. We also observe better scalability w.r.t. both embedding parameters and non-embedding parameters, with GRs leading to 1.5 trillion parameter models, whereas DLRMs performance saturate at about 200 billion parameters.
Finally, all of our main metrics, including Hit Rate@100 and Hit Rate@500 for retrieval, and NE for ranking, empirically scale as a power law of compute used given appropriate hyperparameters. We observe this phenomenon across three orders of magnitude, up till the largest models we were able to test (8,192 sequence length, 1,024 embedding dimension, 24 layers of HSTU), at which point the total amount of compute we used (normalized over 365 days as we use a standard streaming training setting) is close to the total training compute used by GPT-3 (Brown et al., 2020) and LLaMa-2 (Touvron et al., 2023b), as shown in Figure 1. Within a reasonable range, the exact model hyperparameters play less important roles compared to the total amount of training compute applied. In contrast to language modeling (Kaplan et al., 2020), sequence length play a significantly more important role in GRs, and it’s important to scale up sequence length and other parameters in tandem. This is perhaps the most important advantage of our proposed method, as we’ve shown for the first time that scaling law from LLMs may also apply to large-scale recommendation systems.
@@ 5 Related Work
Prior work on sequential recommenders reduces user interactions to a single homogeneous sequence over items (Hidasi et al., 2016; Kang & McAuley, 2018). Industrial-scale applications of sequential approaches are primarily pairwise attention (Zhou et al., 2018) or sequential encoders as part of DLRMs (Chen et al., 2019; Xia et al., 2023). Multi-stage attention has been explored in lieu of self-attention to improve efficiency (Chang et al., 2023). Generative approaches that represent ids as a token series have been explored in retrieval (Zhuo et al., 2020). We give a more extensive discussion of prior work in Section B.1.
@@ Acknowledgements
This work represents the joint efforts of hundreds of people, and would not be possible without work from the following contributors (alphabetical order): Adnan Akhundov, Bugra Akyildiz, Shabab Ayub, Alex Bao, Renqin Cai, Jennifer Cao, Guoqiang Jerry Chen, Lei Chen, Sean Chen, Xianjie Chen, Huihui Cheng, Weiwei Chu, Ted Cui, Shiyan Deng, Nimit Desai, Fei Ding, Francois Fagan, Lu Fang, Liang Guo, Liz Guo, Jeevan Gyawali, Yuchen Hao, Daisy Shi He, Samuel Hsia, Jie Hua, Yanzun Huang, Hongyi Jia, Rui Jian, Jian Jin, Rahul Kindi, Changkyu Kim, Yejin Lee, Fu Li, Hong Li, Shen Li, Wei Li, Zhijing Li, Xueting Liao, Emma Lin, Hao Lin, Jingzhou Liu, Xingyu Liu, Kai Londenberg, Liang Luo, Linjian Ma, Matt Ma, Yun Mao, Bert Maher, Matthew Murphy, Satish Nadathur, Min Ni, Jongsoo Park, Jing Qian, Lijing Qin, Alex Singh, Timothy Shi, Dennis van der Staay, Xiao Sun, Colin Taylor, Shin-Yeh Tsai, Rohan Varma, Omkar Vichare, Alyssa Wang, Pengchao Wang, Shengzhi Wang, Wenting Wang, Xiaolong Wang, Zhiyong Wang, Wei Wei, Bin Wen, Carole-Jean Wu, Eric Xu, Bi Xue, Zheng Yan, Chao Yang, Junjie Yang, Zimeng Yang, Chunxing Yin, Daniel Yin, Yiling You, Keke Zhai, Yanli Zhao, Zhuoran Zhao, Hui Zhang, Jingjing Zhang, Lu Zhang, Lujia Zhang, Na Zhang, Rui Zhang, Xiong Zhang, Ying Zhang, Zhiyun Zhang, Charles Zheng, Erheng Zhong, Xin Zhuang. We would like to thank Shikha Kapoor, Rex Cheung, Lana Dam, Ram Ramanathan, Nipun Mathur, Bo Feng, Yanhong Wu, Zhaohui Guo, Hongjie Bai, Wen-Yun Yang, Zellux Wang, Arun Singh, Bruce Deng, Yisong Song, Haotian Wu, Meihong Wang for product support, and Joseph Laria, Akshay Hegde, Abha Jain, Raj Ganapathy for assistance with program management. Finally, we would like to thank Ajit Mathews, Shilin Ding, Hong Yan, Lars Backstrom for their leadership support, and insightful discussions with Andrew Tulloch, Liang Xiong, Kaushik Veeraraghavan, and Gaofeng Zhao.
@@ Appendix A Notations
We summarize key notations used in this paper in Table 8 and Table 9.
@@
$\Psi_{k}(t_{j})$ The $k$-th training example ($k$ is ordered globally) emitted by the feature logging system at time $t_{j}$. In a typical DLRM recommendation system, after the user consumes some content $\Phi_{i}$ (by responding with an action $a_{i}$ such as skip, video completion and share), the feature logging system joins the tuple $(\Phi_{i},a_{i})$ with the features used to rank $\Phi_{i}$, and emits $(\Phi_{i},a_{i},\text{features for }\Phi_{i})$ as a training example $\Psi_{k}(t_{j})$. As discussed in Section 2.3, DLRMs and GRs deal with different numbers of training examples, with the number of examples in GRs typically being 1-2 orders of magnitude smaller.
$n_{c}$ ($n_{c,i})$ Number of contents that user has interacted with (of user/sample $i$).
$\Phi_{0},\ldots,\Phi_{n_{c}-1}$ List of contents that a user has interacted with, in the context of a recommendation system.
$a_{0},\ldots,a_{n_{c}-1}$ List of user actions corresponding to $\Phi_{i}$s. When all predicted events are binary, each action can be considered a multi-hot vector over (atomic) events such as like, share, comment, image view, video initialization, video completion, hide, etc.
$E,F$ Categorical features in DLRMs, in Figure 2. $E_{0},E_{1}$, $\ldots$, $E_{7},E_{8}$, and $F_{0},F_{1}$, $\ldots$, $F_{7}$ represent transformations of $(\Phi_{0},a_{0},t_{0}),\ldots,(\Phi_{n_{c}-1},a_{n_{c}-1},t_{n_{c}-1})$ obtained at various points in time via feature extraction (e.g., most recent 10 liked images, most similar 50 urls that the user clicked on in the past compared to the current candidate, etc.). “merge & sequentialize” denotes the (virtual) reverse process of obtaining the raw engagement series $(\Phi_{0},a_{0},t_{0}),\ldots,(\Phi_{n_{c}-1},a_{n_{c}-1},t_{n_{c}-1})$.
$G,H$ Categorical features in DLRMs, in Figure 2 that are not related to user-content engagements. These features (e.g., demographics or followed creators) are merged into the main time series (list of contents user engaged with, e.g., $\Phi_{0},a_{0},\ldots,\Phi_{n_{c}-1},a_{n_{c}-1}$), as discussed in Section 2.1 and illustrated in Figure 2.
$n$ ($n_{i}$) Number of tokens in the sequential transduction task (of user/sample $i$). While $O(n)=O(n_{c})$, $n$ can differ from $n_{c}$ even without any non-interaction related categorical features; see e.g., Table 1.
$x_{0},\ldots,x_{n-1}$ List of input tokens in the sequential transduction task.
$y_{0},\ldots,y_{n-1}$ List of output tokens in the sequential transduction task.
@@
$\mathbb{X}$, $\mathbb{X}_{c}$ Vocabulary of all input/output tokens ($\mathbb{X}$) and its content subset $(\mathbb{X}_{c}$).
$N$, $N_{c}$ $\max_{i}n_{i}$, $\max_{i}n_{c,i}$.
@@
$s_{u}(n_{i})$, $\hat{s}_{u}(n_{i})$ Sampling rate for user $i$, used in generative training (Section 2.3).
@@
$d_{qk}$ Attention dimension size in HSTU and Transformers. This applies to $Q(X)$ and $K(X)$ in Equation 1.
$d_{v}$ Value dimension size in HSTU. For Transformers, we typically have $d_{qk}=d_{v}$.
$d_{ff}$ Hidden dimension size in pointwise feedforward layers of Transformers. HSTU does not utilize feedforward layers; see $U(X)$ below.
@@
Table 8: Table of Notations (continued on the next page).
Symbol Description
$X$ Input to an HSTU layer. In standard terminology (before batching), $X\in\mathbb{R}^{N\times d}$ assuming we have a input sequence containing $N$ tokens.
$Q(X),K(X),V(X)$ Query, key, value in HSTU obtained for a given input $X$ based on Equation 1. The definition is similar to $Q$, $K$, and $V$ in standard Transformers. $Q(X),K(X)\in\mathbb{R}^{h\times N\times d_{qk}}$, and $V(X)\in\mathbb{R}^{h\times N\times d_{v}}$.
$U(X)$ HSTU uses $U(X)$ to “gate” attention-pooled values ($V(X)$) in Equation 3, which together with $f_{2}(\cdot)$, enables HSTU to avoid feedforward layers altogether. $U(X)\in\mathbb{R}^{h\times N\times d_{v}}$.
$A(X)$ Attention tensor obtained for input $X$. $A(X)\in\mathbb{R}^{h\times N\times N}$.
$Y(X)$ Output of a HSTU layer obtained for the input $X$. $Y(X)\in\mathbb{R}^{d}$.
$\text{Split}(\cdot)$ The operation that splits a tensor into chunks. $\phi_{1}(f_{1}(X)))\in\mathbb{R}^{N\times(2hd_{qk}+2hd_{v})}$ in Equation 1; we obtain $U(X)$, $V(X)$ (both of shape $h\times N\times d_{v}$), $Q(X)$, $K(X)$ (both of shape $h\times N\times d_{qk}$) by splitting the larger tensor (and permuting dimensions) with $U(X),V(X),Q(X),K(X)=\text{Split}(\phi_{1}(f_{1}(X)))$.
$\text{rab}^{p,t}$ relative attention bias that incorporates both positional (Raffel et al., 2020) and temporal information (based on the time when the tokens are observed, $t_{0},\ldots,t_{n-1}$; one possible implementation is to apply some bucketization function to $(t_{j}-t_{i})$ for $(i,j)$). In practice, we share $\text{rab}^{p,t}$ across different attention heads within a layer, hence $\text{rab}^{p,t}\in\mathbb{R}^{1\times N\times N}$.
@@
$\alpha$ Parameter controlling sparsity in the Stochastic Length algorithm used in HSTU (Section 3.2).
$R$ Register size on GPUs, in the context of the HSTU algorithm discussed in Section 3.2.
$b_{m}$ Microbatch size, in the M-FALCON algorithm discussed in Section 3.4.
Table 9: Table of Notations (continued)
Appendix B Generative Recommenders: Background and Formulations
Many readers are likely more familiar with classical Deep Learning Recommendation Models (DLRMs) (Mudigere et al., 2022) given its popularity from YouTube DNN days (Covington et al., 2016) and its widespread usage in every single large online content and e-commerce platform (Cheng et al., 2016; Zhou et al., 2018; Wang et al., 2021; Chang et al., 2023; Xia et al., 2023; Zhai et al., 2023a). DLRMs operate on top of heterogeneous feature spaces using various neural networks including feature interaction modules (Guo et al., 2017; Xiao et al., 2017; Wang et al., 2021), sequential pooling or target-aware pairwise attention modules (Hidasi et al., 2016; Zhou et al., 2018; Chang et al., 2023) and advanced multi-expert multi-task modules (Ma et al., 2018; Tang et al., 2020). We hence provided an overview of Generative Recommenders (GRs) by contrasting them with classical DLRMs explicitly in Section 2 and Section 3. In this section, we give the readers an alternative perspective starting from the classical sequential recommender literature.
B.1 Background: Sequential Recommendations in Academia and Industry
B.1.1 Academic Research (Traditional Sequential Recommender Settings)
Recurrent neural networks (RNNs) were first applied to recommendation scenarios in GRU4Rec (Hidasi et al., 2016). Hidasi et al. (2016) considered Gated Recurrent Units (GRUs) and applied them over two datasets, RecSys Challenge 2015  [^2] http://2015.recsyschallenge.com/ and VIDEO (a proprietary dataset). In both cases, only positive events (clicked e-commerce items or videos where users spent at least a certain amount of time watching) were kept as part of the input sequence. We further observe that in a classical industrial-scale two-stage recommendation system setup consisting of retrieval and ranking stages (Covington et al., 2016), the task that Hidasi et al. (2016) solved primarily maps to the retrieval task.
Transformers, sequential transduction architectures, and their variants. Advances in sequential transduction architectures in later years, in particular Transformers (Vaswani et al., 2017), have motivated similar advancements in recommendation systems. SASRec (Kang & McAuley, 2018) first applied Transformers in an autoregressive setting. They considered the presence of a review or rating as positive feedback, thereby converting classical datasets like Amazon Reviews  [^3] https://jmcauley.ucsd.edu/data/amazon/ and MovieLens  [^4] https://grouplens.org/datasets/movielens/1m/https://grouplens.org/datasets/movielens/20m/ to sequences of positive items, similar to GRU4Rec. A binary cross entropy loss was employed, where positive target is defined as the next “positive” item (recall this is in essence just presence of a review or rating), and negative target is randomly sampled from the item corpus $\mathbb{X}=\mathbb{X}_{c}$.
Most subsequent research were built upon similar settings as GRU4Rec (Hidasi et al., 2016) and SASRec (Kang & McAuley, 2018) discussed above, such as BERT4Rec (Sun et al., 2019) applying bidirectional encoder setting from BERT (Devlin et al., 2019), S3Rec (Zhou et al., 2020) introducing an explicit pre-training stage, and so on.
B.1.2 Industrial Applications as part of Deep Learning Recommendation Models (DLRMs).
Sequential approaches, including sequential encoders and pairwise attention modules, have been widely applied in industrial settings due to their ability to enhance user representations as part of DLRMs. DLRMs commonly use relatively small sequence lengths, such as 20 in BST (Chen et al., 2019), 1,000 in DIN (Zhou et al., 2018), and 100 in TransAct (Xia et al., 2023). We observe that these are 1-3 orders of magnitude smaller compared with 8,192 in this work (Section 4.3).
The inability to capture longer term user preferences does not represent a problem due to two aspects. First, precomputed user profiles/embeddings (Xia et al., 2023) or external vector stores (Chang et al., 2023) are commonly used in modern DLRMs. Second, a significant number of contextual-, user-, and item-side features were generally employed (Zhou et al., 2018; Chen et al., 2019; Chang et al., 2023; Xia et al., 2023) and various heterogeneous networks, such as FMs (Xiao et al., 2017; Guo et al., 2017), DCNs (Wang et al., 2021), MoEs, etc. are used to transform representations and combine outputs.
In contrast to sequential settings discussed in Section B.1.1, all major industrial work defines loss over (user/request, candidate item) pairs. In the ranking setting, a multi-task binary cross-entropy loss is commonly used. In the retrieval setting, two tower setting (Covington et al., 2016) remains the dominant approach. Recent work has investigated representing the next item to recommend as a probability distribution over a sequence of (sub-)tokens, such as OTM (Zhuo et al., 2020) and DR (Gao et al., 2021) (note that in more recent work, the same setting is sometimes denoted as “generative retrieval”). They commonly utilize beam search to decode the item from sub-tokens. Advanced learned similarity functions, such as mixture-of-logits (Zhai et al., 2023a), have also been proposed and deployed as an alternative to two-tower setting and beam search given proliferation of modern accelerators such as GPUs, custom ASICs, and TPUs.
From a problem formulation perspective, we consider all work discussed above part of DLRMs (Mudigere et al., 2022) given the model architectures, features used, and losses used differ significantly from academic sequential recommender research discussed in Section B.1.1. We’d also like to point out that there have been no reports of successful applications of fully sequential ranking settings in industry, especially not at billion daily active users (DAU) scale, prior to this work.
B.2 Formulations: Ranking and Retrieval as Sequential Transduction Tasks in Generative Recommenders (GRs)
We next discuss three limitations in the traditional sequential recommender settings and DLRM settings, and how Generative Recommenders (GRs) address them from a problem formulation perspective.
Ignorance of features other than user-interacted items. Past sequential formulations only consider contents (items) users explicitly interacted with (Hidasi et al., 2016; Kang & McAuley, 2018; Sun et al., 2019; Zhou et al., 2020), while industry-scale recommendation systems prior to GRs are trained over a vast number of features to enhance the representation of users and contents (Covington et al., 2016; Cheng et al., 2016; Zhou et al., 2018; Chen et al., 2019; Chang et al., 2023; Xia et al., 2023; Zhai et al., 2023a). GR addresses this limitation by a) compressing other categorical features and merging them with the main time series, and b) capturing numerical features through cross-attention interaction utilizing a target-aware formulation as discussed in Section 2.1 and Figure 2. We validate this by showing that the traditional “interaction-only” formulation that ignores such features degrades model quality significantly; experiment results can be found in the rows labeled “GR (interactions only)” in Table 7 and Table 6, where we show utilizing only interaction history led to a 1.3% decrease in hit rate@100 for retrieval and a 2.6% NE decrease in ranking (recall a 0.1% change in NE is significant, as discussed in Sections 4.1.2 and 4.3.1).
User representations are computed in a target-independent setting. A second issue is most traditional sequential recommenders, including GRU4Rec (Hidasi et al., 2016), SASRec (Kang & McAuley, 2018), BERT4Rec (Sun et al., 2019), S3Rec (Zhou et al., 2020), etc. are formulated in a target-independent fashion where for a target item $\Phi_{i}$, $\Phi_{0},\Phi_{1},\ldots,\Phi_{i-1}$ are used as encoder input to compute user representations, which is then used to provide predictions. In contrast, most major DLRM approaches used in industrial settings formulated the sequential modules used in a target-aware fashion, with the ability to incorporate “target” (ranking candidate) information into the user representations. These include DIN (Zhou et al., 2018) (Alibaba), BST (Chen et al., 2019) (Alibaba), TWIN (Chang et al., 2023) (Kwai), and TransAct (Xia et al., 2023) (Pinterest).
Generative Recommenders (GRs) combines the best of both worlds by interleaving the content and action sequences (Section 2.2) to enable applying target-aware attention in causal, autoregressive settings. We categorize and contrast prior work and this work in Table 10  [^5] Most large-scale industrial recommenders need to be trained in a streaming/single-pass setting due to vast amount of logged data..
Input for target item $i$ Expected output for target item $i$ Architecture Training Procedure
GRs $\Phi_{0},a_{0},\Phi_{1},a_{1},\ldots,\Phi_{i}$ $a_{i}$ (target-aware) Self-attention (HSTU) Causal autoregressive (streaming/single-pass)
GRU4Rec $\Phi_{0},\Phi_{1},\ldots,\Phi_{i-1}$ $\Phi_{i}$ RNNs (GRUs) Causal autoregressive (multi-pass)
SASRec Self-attention (Transformers)
BERT4Rec $\Phi_{0},\Phi_{1},\ldots,\Phi_{i-1}$ (at inference time) $\Phi_{i}$ Self-attention (Transformers) Sequential multi-pass  [^6] BERT4Rec leverages multi-pass training with a mixture of Cloze and pointwise (last item) supervision losses; S3Rec utilizes multi-pass training with pre-training and finetuning as two separate stages.
S3Rec
DIN $\Phi_{0},\Phi_{1},\ldots,\Phi_{i}$ $a_{i}$ (target aware, implicitly as part of DLRMs) Pairwise attention Pointwise (generally streaming/single pass)
BST Self-attention (Transformers)
TWIN Two-stage pairwise attention
TransAct $(\Phi_{0},a_{0}),\ldots,(\Phi_{i-1},a_{i-1}),\Phi_{i}$ Self-attention (Transformers)
Table 10: Comparison of prior work on sequential recommenders and GRs, in the ranking setting, with DLRMs included for completeness.
Discriminative formulations restrict applicability of prior sequential recommender work to pointwise settings. Finally, traditional sequential recommenders are discriminative by design. Existing sequential recommender literature, including seminal work such as GRU4Rec and SASRec, model $p(\Phi_{i}|\Phi_{0},a_{0},\ldots,\Phi_{i-1},a_{i-1})$, or the conditional distribution of the next item to recommend given users’ current states. On the other hand, we observe that there are two probabilistic processes in standard recommendation systems, namely the process of the recommendation system suggesting a content $\Phi_{i}$ (e.g., some photo or video) to the user, and the process of the user reacting to the suggested content $\Phi_{i}$ via some action $a_{i}$ (which can be a combination of like, video completion, skip, etc.).
A generative approach needs to model the joint distribution over the sequence of suggested contents and user actions, or $p(\Phi_{0},a_{0},\Phi_{1},a_{1},\ldots,\Phi_{n_{c}-1},a_{n_{c}-1})$, as discussed in Section 2.2. Our proposal of Generative Recommenders enables modeling of such distributions, as shown in Table 11 (Figure 8). Note that the next action token ($a_{i}$) prediction task is exactly the GR ranking setting discussed in Table 1, whereas the next content ($\Phi_{i}$) prediction task is similar to the retrieval setting adapted to the interleaved setting, with the target changed in order to learn the input data distribution.
Task Specification (Inputs / Outputs / Length)
Next action token ($a_{i}$) prediction $x_{i}$s $\Phi_{0},a_{0},\Phi_{1},a_{1},\ldots,\Phi_{n_{c}-2},a_{n_{c}-2},\Phi_{n_{c}-1},a_{n_{c}-1}$
$y_{i}$s $a_{0},\varnothing,a_{1},\varnothing,\ldots,a_{n_{c}-2},\varnothing,a_{n_{c}-1},\varnothing$
$n$ $2n_{c}$
Next content token ($\Phi_{i}$) prediction $x_{i}$s $\Phi_{0},a_{0},\Phi_{1},a_{1},\ldots,\Phi_{n_{c}-2},a_{n_{c}-2},\Phi_{n_{c}-1},a_{n_{c}-1}$
$y_{i}$s $\varnothing,\Phi_{1},\varnothing,\Phi_{2},\ldots,\varnothing,\Phi_{n_{c}-1},\varnothing,\varnothing$
$n$ $2n_{c}$
Table 11: Generative modeling over $p(\Phi_{0},a_{0},\ldots,\Phi_{n_{c}-1},a_{n_{c}-1})$. An illustration is provided in Figure 8.
Refer to captionRefer to caption
Figure 8: Comparison of traditional sequential recommenders (left) and Generative Recommenders (right). We illustrate sequential recommenders in causal autoregressive settings and GRs without contextual features to facilitate comparison. On the left hand side, the action types $a_{i}$s are either ignored or combined with item information $\Phi_{i}$s using MLPs, before going into self-attention blocks.
Importantly, this formulation not only enables proper modeling of data distribution but further enables sampling sequences of items to recommend to the user directly via e.g., beam search. We hypothesize that this will lead to a superior approach compared with traditional listwise settings (e.g., DPP (Gillenwater et al., 2014) and RL (Zhao et al., 2018)), and we leave the full formulation and evaluation of such systems (briefly discussed in Section 6) as a future work.
Appendix C Evaluation: Synthetic Data